OpenStudy (anonymous):

Convert the decimal to a percent. 3.45 A. 345% B. 34.5% C. 3.45% D. 0.345%

OpenStudy (anonymous):

@Nurali @Destinymasha

OpenStudy (anonymous):

To convert a decimal into a percent, you simply need to multiply said decimal by 100. So, 3.45 * 100 = 345% !!

OpenStudy (mathstudent55):

Since percent means "divide by 100," changing from percent to a decimal means "multiply by 100."

OpenStudy (anonymous):

Convert the percent to a decimal. 12.5% A. 0.00125 B. 0.125 C. 1.25 D. 12.5

OpenStudy (anonymous):

With this question, you do the opposite of what you did before. Take the decimal and divide by 100. 12.5 / 100 = .125

OpenStudy (anonymous):

Which percent is equal to the given fraction? A. 3.5% B. 30% C. 35% D. 60%

OpenStudy (nurali):

3/5=0.6 0.6*100=60%